# Transmission Mask Analysis for Range–Doppler Sensing in Half-Duplex ISAC

### v1

@@ II-B Performance Metrics
    Moderately dynamic regime $\nu<M$: The Doppler-induced complex exponential accumulates at most one full $2\pi$ phase rotation within a single PRI. In this case, we may first perform range compression within each PRI. Doppler frequencies are then estimated by applying spectral estimators to the sequence of range-compressed samples across pulse repetition interval (PRI), where the sampling rate is the PRF.
@@
    Highly dynamic regime $\nu\geqslant M$: In this case, the range compression and Doppler estimation operations can no longer be decoupled.
@@ III Analysis of the Range-Doppler Response
In this section, we analyze the characteristics of $\mathbb{E}\{|r(k,l,\nu)|^{2}\}$.
@@ III-A The range sidelobes $k\neq l$
which turns out to be independent of the Doppler shift $\nu$. Moreover, since $m_{\rm t}[n]$ and $m_{\rm r}[n]$ both have period $N$, we further have
@@ III-C The range mainlobes $k=l$, Highly Dynamic Regime
Another issue that raises concern is the mainlobe fluctuation phenomenon, a unique characteristic of half-duplex sensing schemes, which refers to the unevenness of $\mathbb{E}\{|r(k,k,0)|^{2}$ at different values of $k$. Since there is no Doppler shift in this case, we may directly apply the result in [15]:
@@ IV Numerical Results
Figure 1: The range-Doppler responses of different masks at range mainlobes ($k=l$), full Doppler span $\nu\in\{0,\ldots,MN-1\}$.
@@ IV-1 Range mainlobe ($k=l$), global Doppler behavior (highly dynamic regime)
Fig. 1 plots $\mathbb{E}\{|r(k,k,\nu)|^{2}\}$ over the full Doppler span $\nu\in\{0,\ldots,MN-1\}$. The periodic masking structure manifests itself as a pronounced, highly structured ambiguity pattern. In particular, the comb-like achieves a considerably low Doppler sidelobe level at most Doppler bins, at the cost of a grating-lobe-like behavior at certain Doppler bins, due to its strict periodicity. By contrast, the Singer CDS and random masks yield markedly different global distributions. While both exhibit ambiguity induced by periodic operation, the CDS tends to provide a more favorable global range-Doppler profile as discussed in Sec. III (e.g., worst-case behavior along the range mainlobe), whereas the random mask shows less controlled variations. We note that the comb-like mask has a smaller $\rho$. According to (5), its sidelobe levels would be higher if it also had $\rho=31/63$. Consequently, as portrayed in Fig. 1d, the worst case mean Doppler sidelobe level over $k$ of the comb-like mask become larger than that of the CDS, after rescaling by $\rho$.
@@ IV-2 Range mainlobe ($k=l$), zoomed in (moderately dynamic regime)
Fig. 2 zooms into the first Doppler block $\nu\in\{0,\ldots,M-1\}$ of Fig. 1, corresponding to the moderately dynamic regime. In this regime, as discussed in Sec. III, the deterministic Doppler-induced term $|S_{k,MN}(\nu)|^{2}$ does not contribute to nonzero Doppler bins within this local window, and the remaining behavior is controlled by the mask-dependent factor $a[k]$. Accordingly, masks that equalize $a[k]$ yield smaller and more stable local Doppler sidelobes. The Singer CDS mask exhibits the most favorable local behavior, which is consistent with its minimax optimality in the moderately dynamic regime, while the random mask shows larger mainlobe fluctuation and higher sidelobe levels. The comb-like mask, although does not exhibit the grating lobes in this local window, suffers from the most severe mainlobe fluctuation.
@@
IV-A Range-Doppler response at a fixed $k$
Fig. 3 examines the range-Doppler response at a specific true range bin $k=20$, with varying trial range bin $l$. The two 2D plots show that the sidelobe surface is essentially flat along the Doppler axis, supporting the Doppler-invariance property for $k\neq l$ established in Sec. III. Consequently, one should evaluate the sensing performance of different masks in the range-sidelobe domain by considering their range-domain correlation structures. This is further highlighted by the zero-Doppler slice portrayed in Fig. 3c, where the Singer CDS exhibits the expected sidelobe control advantage over both the random mask and the comb-like mask.
@@ IV-A Range-Doppler response at a fixed $k$
(c) Zero-Doppler slice

### v2

@@ II-B Performance Metrics
    Moderately dynamic regime ($\nu_{0}<M$): In this regime, all targets have true Doppler shifts $\nu_{0}$ bounded by $M$. Note that $\nu$ is defined as the trial Doppler minus the true Doppler ($\nu=\nu_{\rm t}-\nu_{0}$). Under the condition $\nu_{0}<M$ for all targets, we only need to consider $\mathbb{E}\{|r(k,l,\nu)|^{2}\}$ in the region $\nu<M$, since larger Doppler mismatches will not occur. In practice, this means range and Doppler processing can be decoupled [^1] In terms of performance, this treatment works well only when $\nu_{0}\ll M$.: one can first perform range compression within each PRI, and then estimate Doppler frequencies by applying spectral estimation across the sequence of range-compressed samples, with sampling rate equal to the PRF.
@@
    Highly dynamic regime ($\nu_{0}\geq M$): If any target has a true Doppler $\nu_{0}$ exceeding $M$, the above separation no longer holds, necessitating joint range-Doppler processing. In this case, we would have to take into account $\mathbb{E}\{|r(k,l,\nu)|^{2}\}$ in the region $\nu\geqslant M$.
@@ III Analysis of the Range-Doppler Response
In this section, we analyze $\mathbb{E}\{|r(k,l,\nu)|^{2}\}$. In particular, we only consider Doppler differences at $\nu=\{0,1,\dotsc,MN-1\}$. The characteristic at fractional Doppler differences can be adjusted using the windowing technique, and can be depicted using the “iceberg–sea-level” decomposition [8].
@@ III-A The range sidelobes $k\neq l$
which turns out to be independent of the Doppler difference $\nu$. Moreover, since $m_{\rm t}[n]$ and $m_{\rm r}[n]$ both have period $N$, we further have
@@ III-C The range mainlobes $k=l$, Highly Dynamic Regime
Another issue that raises concern is the mainlobe fluctuation phenomenon, a unique characteristic of half-duplex sensing schemes, which refers to the unevenness of $\mathbb{E}\{|r(k,k,0)|^{2}$ at different values of $k$. Since there is no Doppler difference in this case, we may directly apply the result in [16]:
@@
Figure 1: The range-Doppler responses of different masks at range mainlobes ($k=l$), full Doppler difference span $\nu\in\{0,\ldots,MN-1\}$.
IV Numerical Results
@@ IV-1 Range mainlobe ($k=l$), global Doppler behavior (highly dynamic regime)
Fig. 1 plots $\mathbb{E}\{|r(k,k,\nu)|^{2}\}$ over the full Doppler difference span $\nu\in\{0,\ldots,MN-1\}$. The periodic masking structure manifests itself as a pronounced, highly structured ambiguity pattern. In particular, the comb-like achieves a considerably low Doppler sidelobe level at most Doppler bins, at the cost of a grating-lobe-like behavior at certain Doppler bins, due to its strict periodicity. By contrast, the Singer CDS and random masks yield markedly different global distributions. While both exhibit ambiguity induced by periodic operation, the CDS tends to provide a more favorable global range-Doppler profile as discussed in Sec. III (e.g., worst-case behavior along the range mainlobe), whereas the random mask shows less controlled variations. We note that the comb-like mask has a smaller $\rho$. According to (5), its sidelobe levels would be higher if it also had $\rho=31/63$. Consequently, as portrayed in Fig. 1d, the worst case mean Doppler sidelobe level over $k$ of the comb-like mask become larger than that of the CDS, after rescaling by $\rho$.
@@ IV-2 Range mainlobe ($k=l$), zoomed in (moderately dynamic regime)
Fig. 2 zooms into the first Doppler difference block $\nu\in\{0,\ldots,M-1\}$ of Fig. 1, corresponding to the moderately dynamic regime. In this regime, as discussed in Sec. III, the deterministic Doppler-induced term $|S_{k,MN}(\nu)|^{2}$ does not contribute to nonzero Doppler bins within this local window, and the remaining behavior is controlled by the mask-dependent factor $a[k]$. Accordingly, masks that equalize $a[k]$ yield smaller and more stable local Doppler sidelobes. The Singer CDS mask exhibits the most favorable local behavior, which is consistent with its minimax optimality in the moderately dynamic regime, while the random mask shows larger mainlobe fluctuation and higher sidelobe levels. The comb-like mask, although does not exhibit the grating lobes in this local window, suffers from the most severe mainlobe fluctuation.
@@
(c) Zero-Doppler-difference slice
@@
IV-A Range-Doppler response at a fixed $k$
Fig. 3 examines the range-Doppler response at a specific true range bin $k=20$, with varying trial range bin $l$. The two 2D plots show that the sidelobe surface is essentially flat along the Doppler axis, supporting the Doppler-invariance property for $k\neq l$ established in Sec. III. Consequently, one should evaluate the sensing performance of different masks in the range-sidelobe domain by considering their range-domain correlation structures. This is further highlighted by the zero-Doppler-difference slice portrayed in Fig. 3c, where the Singer CDS exhibits the expected sidelobe control advantage over both the random mask and the comb-like mask.